### v1

# TUS-REC2024: A Challenge to Reconstruct 3D Freehand Ultrasound Without External Tracker

Abstract
Trackerless freehand ultrasound reconstruction aims to reconstruct 3D volumes from sequences of 2D ultrasound images without relying on external tracking systems, offering a low-cost, portable, and widely deployable alternative for volumetric imaging. However, it presents significant challenges, including accurate inter-frame motion estimation, minimisation of drift accumulation over long sequences, and generalisability across scanning protocols. The TUS-REC2024 Challenge was established to benchmark and accelerate progress in trackerless 3D ultrasound reconstruction by providing a publicly available dataset for the first time, along with a baseline model and evaluation framework. The Challenge attracted over 43 registered teams, of which 6 teams submitted 21 valid dockerized solutions. Submitted methods spanned a wide range of algorithmic approaches, including recurrent models, registration-driven volume refinement, attention, and physics-informed models. This paper presents an overview of the Challenge design, summarises the key characteristics of the dataset, provides a concise literature review, introduces the technical details of the underlying methodology working with tracked freehand ultrasound data, and offers a comparative analysis of submitted methods across multiple evaluation metrics. The results highlight both the progress and current limitations of state-of-the-art approaches in this domain, and inform directions for future research. The data, evaluation code, and baseline are publicly available to facilitate ongoing development and reproducibility. As a live and evolving benchmark, this Challenge is designed to be continuously developed and improved. The Challenge was held at MICCAI 2024 and will be organised again at MICCAI 2025, reflecting its growing impact and the sustained commitment to advancing this field.
keywords:
Trackerless , Freehand , Ultrasound , 3D reconstruction , Spatial transformation estimation , MICCAI 2024 challenge , TUS-REC2024
[…]
[ucl] organization=UCL Hawkes Institute, Department of Medical Physics and Biomedical Engineering, University College London,city=London, postcode=WC1E 6BT, country=U.K.
[…]
[MUSIC_Lab_1] organization=National-Regional Key Technology Engineering Laboratory for Medical Ultrasound, School of Biomedical Engineering, Health Science Center, Shenzhen University,city=Shenzhen, postcode=518000, country=China
[…]
[Madison_Elastography_Lab_1] organization=Department of Medical Physics, University of Wisconsin (UW) School of Medicine and Public Health, city=Madison, postcode=WI 53705, country=USA
[…]
[kcl] organization=School of Biomedical Engineering & Imaging Sciences, King’s College London,city=London, postcode=WC2R 2LS, country=U.K.
1 Introduction
Ultrasound imaging remains a cost-effective, non-invasive modality with real-time capabilities, making it a valuable tool across a wide range of clinical applications. Although the underlying data is inherently three-dimensional (3D), standard ultrasound typically captures single two-dimensional (2D) frame without spatial localisation across frames. This poses challenges for applications requiring accurate volumetric information, such as biometric quantification, image registration, and 3D visualisation. While expert clinicians can often infer 3D structure mentally or through standardised acquisition protocols (e.g., standard planes), the absence of inter-frame positional data limits reproducibility and the integration of ultrasound into advanced image analysis workflows.
Ongoing work seeks to address this limitation by using 3D ultrasound probes to enable 3D reconstruction. 3D ultrasound probes is capable of acquiring volumetric data directly, using dedicated mechanical probe or 2D array transducers. While these probes provide valuable 3D imaging capabilities and offer flexible scanning trajectories, their higher cost and limited availability may restrict their use in some clinical settings, such as low-resource environments, point-of-care scenarios, or mobile and emergency units where portability and affordability are critical.
A key advantage of freehand 2D ultrasound is its widespread availability and long-standing integration into clinical workflows. It has been used for decades across a broad range of applications, and clinicians are highly familiar with both its handling and interpretation. Building on this established foundation, tracker-based freehand ultrasound reconstruction techniques have been introduced to enable the generation of 3D anatomical representations. These methods aim to enhance conventional 2D ultrasound by incorporating spatial information from external tracking systems, such as optical or electromagnetic (EM) trackers. This enables conventional 2D ultrasound probes to be used for 3D imaging, providing a more flexible and accessible solution in clinical and research applications where dedicated and bulky 3D ultrasound systems may be impractical. However, optical and EM tracking systems present additional challenges in clinical environments. Optical tracking requires a continuous, unobstructed line of sight between the tracker and the camera, although some solutions, such as using multiple cameras, have been proposed to mitigate this limitation. EM tracking remains sensitive to nearby metal objects and electromagnetic interference, which can affect accuracy.
Trackerless freehand ultrasound reconstruction refers to a class of techniques aimed at generating 3D volumetric representations from sequential 2D ultrasound frames, without the use of external tracking systems. Instead, these methods compute the relative spatial transformations among frames using internal data sources, such as image contents and signals from internal sensors. Common approaches include image-based registration, speckle decorrelation, inertial measurement unit (IMU) integration, or learning-based motion estimation such as convolutional neural networks or recurrent models. Additionally, trackerless freehand ultrasound reconstruction may further enhance existing 3D ultrasound systems, rather than serving solely as alternatives.
However, trackless reconstruction remains challenging due to: 1) the difficulty of maintaining accuracy over long sequences of ultrasound frames, where small frame-to-frame errors can accumulate significantly; and 2) the high variability across different tasks and datasets, which complicates the validation and fair comparison of methods. While benchmarking is essential to address this variability, progress has been limited by the scarcity of publicly available datasets, which are critical for both performance evaluation and the development of learning-based approaches.
This application involves practical challenges such as handling both 2D and 3D imaging data, incorporating tracking information, and managing multiple spatial coordinate systems, all of which contribute to a significant barrier for newcomers to the field and may impede broader progress and adoption. Furthermore, comparisons of methods in the existing literature are often conducted on relatively small, private datasets, using a variety of evaluation metrics to assess performance. This variability complicates the comparison of strengths and weaknesses across methods and may lead to biased conclusions, due to dataset characteristics, evaluation metric choices, and inherent differences in the methods’ underlying assumptions. For example, learning-based approaches may assume that training and testing data come from similar distributions, while classical methods may rely on consistently available speckle patterns.
To address these gaps, we present a significant study of trackerless freehand ultrasound reconstruction, aiming to provide researchers with clear technical insights and consistent terminology. This effort is formalised in the form of the TUS-REC2024 Challenge, which is designed to foster both algorithmic innovation and practical clinical applicability by promoting reproducibility, benchmarking, and methodological transparency. TUS-REC2024 Challenge provides a large-scale in vivo ultrasound dataset, consisting of scans from both the left and right forearms of 85 volunteers (2,040 scans, 1,025,448 frames in total), acquired using a time-synchronised optical tracking system. We aim to conduct a comprehensive comparison among methods, evaluating their strengths and weaknesses on a common, large-scale dataset, using a consistent set of carefully-defined performance metrics. This approach will ensure a more objective and transparent assessment of the methods’ relative efficacy, and more importantly, to drive the development of new techniques for freehand ultrasound reconstruction.
The significance of this Challenge lies in three key contributions. First, it establishes a rigorous and standardised benchmarking framework for trackerless freehand ultrasound reconstruction, advancing the development of novel algorithms and promoting objective performance evaluation through withheld test data and unified assessment metrics. Second, it provides the necessary infrastructure to support this benchmarking effort, including the public release of a comprehensive freehand ultrasound dataset, the largest publicly available dataset to date in the field, alongside detailed preliminary materials and accompanying code that describe the end-to-end pipeline for trackerless reconstruction. Third, beyond the outcomes of the Challenge itself, this summary paper delivers additional insights, including a detailed comparative analysis of the participating algorithms and a discussion of the design choices and performance trade-offs for future method development.
Sections 2 and  3 summarise the state of the art methods and provide an overview of trackerless freehand ultrasound reconstruction. Section 4 details the parameters of the Challenge, including the dataset, evaluation metrics and other Challenge setups. Section 5 describes the participation statistics and the methodologies submitted by participating teams, accompanied by the performance analysis of each method. Section 6 discusses the limitations of the Challenge and outlines potential directions for future work. Finally, Section 7 concludes the study by summarising the outcomes of the Challenge and highlighting its key contributions, benefits, and future directions.
2 Related Work
The field of 3D freehand ultrasound reconstruction has undergone a significant evolution over the past two decades. Early trackerless methods relied on statistical modeling to approximate probe motion without external tracking systems, such as speckle decorrelation analysis [1] and linear regression-based motion estimation [2]. While these approaches established the feasibility of trackerless scanning, they were often limited in robustness and generalisability, particularly in the presence of non-linear motion or varying anatomical structures [3, 4].
Recent advances in deep learning have led to a new wave of data-driven methods for trackerless 3D ultrasound reconstruction. Early CNN-based approaches [5] demonstrated that learned image representations could outperform traditional speckle-based techniques in estimating inter-frame motion. Subsequent work expanded on this by incorporating spatial and temporal modeling through Long Short-Term Memory (LSTM) networks [6, 7, 8, 9], and attention [10, 11], enabling more stable trajectory estimation over longer scan sequences. Transformer-based architectures [12, 13] have recently been introduced to better model long-range dependencies and spatial coherence.
Recent research in freehand 3D ultrasound reconstruction has pursued multiple directions to improve motion estimation accuracy, reconstruction speed, and generalisability across protocols and anatomies. One of the promising direction is the fusion of deep learning-based trajectory estimation with volumetric consistency optimisation. For instance, several methods combine learning-based inter-frame pose estimation with multi-view or global model refinement to improve alignment and robustness across large scan sequences [14, 15]. Another active area focuses on addressing domain shifts introduced by different ultrasound transducers. Domain adaptation strategies have been proposed to improve cross-device generalisation, particularly through the learning of transducer-invariant features [16]. A notable recent development involves the application of implicit neural representations to ultrasound imaging. These methods aim to compress the volumetric information and encode it as parameters of a model, offering potential gains in memory efficiency and spatial resolution compared with explicit representation [17, 18, 19]. [20] further extends this concept by adapting Gaussian Splatting techniques to volumetric ultrasound by replacing projection-based rendering with a model that aligns with ultrasound wave propagation. By leveraging anisotropic Gaussians, this method enables faster, memory efficient 3D ultrasound reconstruction.
Geometric and motion-aware modeling have also emerged as a powerful tool for improving reconstruction accuracy. Recent approaches integrate optical flow and spatiotemporal attention to better capture dense inter-frame motion and global spatiotemporal consistency [21, 22], while point cloud-based registration [23] techniques offer motion-driven alternatives to traditional pose estimation pipelines. In parallel, several works have investigated sensor signals and auxiliary tracking modalities to enhance pose estimation accuracy. IMUs have been integrated with learning-based systems to provide motion information or correct drift in trackerless trajectories [5, 24, 25, 26, 27]. Additionally, an evaluation of low-cost tracking alternatives has shown no statistically significant difference between high and low-end optical trackers, further supporting the feasibility of cost-effective freehand 3D ultrasound setups [28].
Despite these promising developments, several limitations persist. First, prior studies have typically relied on private or small-scale datasets, often with only 12-40 subjects, which limits both the statistical power of the findings and the generalisability of trained models. Second, the absence of standardised datasets and evaluation protocols has made it difficult to perform fair and reproducible comparisons between methods. As a result, reported performance varies widely across studies and may be influenced by dataset-specific biases or tuning.
In addition to the above-discussed practical limitations, technical difficulties are also found in other applied machine learning methods in medical imaging. The core lies in accurately estimating the 3D motion of the ultrasound probe from 2D image sequences alone. Unlike tracker-based systems, trackerless methods must compute probe motion from intensity patterns in the images. However, these patterns can be affected by factors that complicate motion estimation, such as tissue deformation due to varying probe pressure and changes in scanning speed or angle. This leads to accumulated drift over time, particularly in long or complex scanning trajectories, which directly impacts reconstruction accuracy and downstream clinical applications especially those requiring spatially accurate 3D volumes.
To date, few effort has been made on large-scale public benchmarking for trackless freehand ultrasound reconstruction. This Challenge was established to fill this gap by offering a comprehensive platform for evaluating trackerless freehand ultrasound reconstruction methods, providing a fair and rigorous environment for performance comparison. In summary, while significant progress has been made in trackerless ultrasound reconstruction, consistent benchmarking is lacking. This Challenge aims to establish a foundation for reproducible and scalable evaluation for trackerless freehand ultrasound reconstruction.
3 Preliminaries
The goal of freehand ultrasound reconstruction is to estimate the transformation between pairs of ultrasound frames within a scan without relying on any external tracking device, thereby enabling the reconstruction of 2D ultrasound images into 3D space.
3.1 Coordinate Systems and Spatial Transformations
Table 1 summarises the terminologies commonly used in freehand ultrasound reconstruction. In learning-based algorithm development, a tracking system is typically used to directly capture the transformation of each ultrasound frame, providing ground truth transformations for training supervision and for evaluating trackerless algorithms. The most commonly utilised tracking modalities are optical tracking systems and EM tracking systems. The optical tracking system consists of a camera that captures pose information and a rigid tracking tool attached to the ultrasound probe [29]. This tool typically includes at least three passive or active markers, which enable the determination of the probe’s six-degree-of-freedom (6-DoF) pose. After spatial calibration, as detailed in Section 3.2, the system can obtain the transformation matrix corresponding to each ultrasound frame. The tracking data are subsequently transferred and stored using an interface such as the open-source PLUS platform [30]. While ultrasound machine and tracking device typically have their own API for data access, tools such as PLUS simplify the process by offering a unified interface and consistent data and file formats, making practical integration more convenient, though not strictly necessary. The EM tracking system [31] comprises three main components: the transmitter, the system control unit, and the tracked receiver. When the probe is moved within the magnetic field produced by the transmitter, the receiver mounted on the probe detects induced electrical currents. The remainder of this section describes the three coordinate systems involved in freehand ultrasound reconstruction, using an optical tracking system as a representative example.
There are three coordinate systems, as shown in Fig. 1(a): the image coordinate system, the tracker tool coordinate system, and the camera (or world) coordinate system. The image coordinate system defines the positions of pixels (in 2D) or voxels (in 3D) within an image. In the context of freehand ultrasound reconstruction, the image coordinate system specifically refers to the 2D coordinate system of individual ultrasound image frame. The transformation recorded by the optical camera is from tracker tool coordinate system to camera coordinate system, which in turn represents the location of the tracker in camera coordinate system. However, this tracker-reported transformation does not directly provide transformation to the coordinate system of the ultrasound image itself, from the other two coordinate systems. Consequently, a transformation, commonly referred to as the calibration matrix, is necessary to map the ultrasound image coordinates to the tracker tool coordinate system. This transformation is crucial for converting each pixel in the 2D ultrasound image to its corresponding voxel in the reconstructed 3D volume. The calibration matrix defines the transformation between the image coordinate system and the tracker tool coordinate system. It incorporates both a scaling factor that converts image coordinates from pixels to millimeters, as well as the spatial calibration that establishes the transformation between the image coordinate system (in millimeters) and the tracker tool coordinate system (in millimeters).
Table 1: Terminologies in freehand ultrasound reconstruction.
Terminology Definition Example Origin Position Unit Example Axis Directions
Image coordinate system A 2D coordinate system defining pixel positions in an image Top-left corner pixel X axis: along the image width, increasing from left to right; Y axis: along the image height, increasing from top to bottom; Z axis: perpendicular to the image plane, increasing into the image.
Tracker tool coordinate system A 3D coordinate system defined by three or four sphere markers which are attached to a rigid body with a unique geometry Origin of the marker attached to the object of interest (phantom, cadaver, patient, etc.) mm As defined by the tracking system / marker manufacturer
Camera (or world) coordinate system A 3D coordinate system defined by the tracking system manufacturer Origin of the tracking system (midpoint between the two camera lenses) mm X axis: increasing downward from the center between the two lenses; Y axis: increasing toward the camera’s right; Z axis: inward, toward the back of the device.
Terminology Definition
$T$ The transformation between two coordinate systems, which changes the coordinate of the same point represented in one coordinate system to another.
$T_{scale}$ The transformation from the image coordinate system (in pixels) to the image coordinate system (in millimeters).
$T_{rotation}$ The transformation from the image coordinate system (in millimeters) to the tracker tool coordinate system (in millimeters).
$T_{i}^{camera\leftarrow tool}$ The transformation from the tracker tool coordinate system (in millimeters) of frame $i$ to the camera coordinate system (in millimeters).
$T_{j\leftarrow i}^{tool}$ The transformation from the tracker tool coordinate system (in millimeters) of frame $i$ to that of frame $j$.
$T_{j\leftarrow i}$ The transformation from image coordinate system (in millimeters) of frame $i$ to that of frame $j$.
Figure 1: (a): Schematic illustration of three coordinate systems: the image coordinate system, the tracker tool coordinate system, and the camera (or world) coordinate system. (b) Schematic illustration of the calibration setup for freehand ultrasound calibration, where an ultrasound probe with an attached tracker tool images a pinhead submerged in a water bath. The pinhead acts as a calibration target, allowing computation of the spatial transformation between the ultrasound image coordinate system and the tracker tool coordinate system.
Let $T_{i}^{camera\leftarrow tool}$ denote the transformation matrix of frame $i$, recorded by the optical camera, representing the transformation from the tracker tool coordinate system of frame $i$ to the camera coordinate system. Then, the rigid transformation from image coordinate system (in millimeters) of frame $i$ to image coordinate system (in millimeters) of frame $j$, $T_{j\leftarrow i}$, is given by:
$T_{j\leftarrow i}=T_{rotation}^{-1}\cdot T_{j\leftarrow i}^{tool}\cdot T_{rotation}$ (1)
where $T_{j\leftarrow i}^{tool}$ denotes the transformation from the $i^{th}$ tracker tool to the $j^{th}$ tracker tool. $T_{rotation}=\left[\begin{array}[]{cc}\mathbf{R_{3\times 3}}&\mathbf{t_{3\times 1}}\\ \mathbf{0}&1\end{array}\right]$ represents the spatial calibration between the image coordinate system (in millimeters) and the tracker tool coordinate system, where $\mathbf{R}$ is a $3\times 3$ rotation matrix and $\mathbf{t}$ is a $3\times 1$ translation vector. The transformation $T_{rotation}$ is obtained through the calibration process, as described in Section. 3.2, while $T_{j\leftarrow i}^{tool}$ can be computed using Eq. (2).
$T_{j\leftarrow i}^{tool}=(T_{j}^{camera\leftarrow tool})^{-1}\cdot T_{i}^{camera\leftarrow tool}$ (2)
Reconstructing the 3D ultrasound volume and the trajectory of the ultrasound frames require determining the position of each frame. Let the first frame serve as the reference frame. If the transformations from each frame to the reference frame are known, the coordinates of all pixels within the scan can be computed using Eq. (3).
$P_{(x,y,z)}=T_{1\leftarrow i}\cdot T_{scale}\cdot p_{(u,v)}$ (3)
where $p_{(u,v)}=(u,v,0,1)^{T}$ and $P_{(x,y,z)}=(x,y,z,1)^{T}$ represent the coordinates of pixel $(u,v)$ in the image coordinate system of the $i^{th}$ frame (in pixels) and the image coordinate system of the first frame (in millimeters), respectively. $T_{scale}=diag(s_{x},s_{y},1,1)$ is the scaling factor that converts from pixels to millimeters, represented by a diagonal matrix with elements $s_{x}$,$s_{y}$,$1$,$1$ along the main diagonal. $s_{x}$ and $s_{y}$ represent the scaling factor along the $x$ and $y$ axes respectively. The two $1$s represent no scaling on the $z$-axis and for the homogeneous coordinate.
When the distance between two ultrasound frames is too large, predicting the transformation becomes challenging. Therefore, a common approach is to estimate the transformation between two adjacent frames and then accumulate these transformations to obtain the desired result. The transformation from the $i^{th}$ frame to the first frame, $T_{1\leftarrow i}$, can be computed by recursively multiplying the previously estimated relative transformations, as shown in Eq. (4).
$T_{1\leftarrow i}=T_{1\leftarrow 2}\cdot T_{2\leftarrow 3}\cdots T_{i-1\leftarrow i}$ (4)
Moreover, Eq. (4) illustrates that estimation errors can propagate and accumulate along the chain, ultimately leading to trajectory drift. To mitigate this issue, [7] leverages the long-term dependencies within the ultrasound sequence to estimate the transformations between frames at intermediate intervals.
3.2 Calibration
The calibration process in freehand ultrasound reconstruction involves both spatial and temporal components. Temporal calibration ensures synchronisation between the timestamped ultrasound image frames acquired from the ultrasound machine and the corresponding transformation data recorded by the optical tracking system. This calibration can be performed using the PLUS Toolkit [30], as well as other established methods described in the literature [29, 32, 33, 34]. Synchronisation can be achieved by identifying the optimal time offset that maximises correlation between the probe’s motion (from tracking) and the observed motion in the image stream. Spatial calibration is required to determine the transformation between the ultrasound image coordinate system (in pixels) and the tracker tool coordinate system. Since the optical tracker records the pose of the tracker tool relative to the camera coordinate system, the calibration matrix enables accurate mapping of the ultrasound frames into the spatial coordinate system of the tracking environment.
In this study, a pinhead based method was employed for spatial calibration (as shown in Fig. 1(b)). A pinhead served as the calibration phantom and was repeatedly imaged in the ultrasound images while simultaneously recording the corresponding transformation matrices from the optical tracker. During data acquisition, the ultrasound probe, equipped with tracking markers, was moved at various angles and distances relative to the pinhead. Additionally, the pinhead was designed to appear at different locations within the ultrasound image plane. The calibration was performed in a water medium to ensure optimal ultrasound imaging quality.
The goal of the calibration process is to estimate the transformation matrix that transforms 2D points from the image coordinate system to the camera coordinate system. Let $\{p_{(u_{i},v_{i})}|i=1,...,n\}$ denote the set of 2D coordinates of the pinhead in the ultrasound image coordinate system, and $P$ the corresponding 3D coordinates in the camera coordinate system. Since the pinhead remains stationary throughout the acquisition, each 2D image point, when transformed into 3D space using the estimated calibration matrix, should be expected to converge at the same 3D location $P$, as shown in Eq. (3.2).
$P$ $=$ $T_{1}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{1},v_{1})}$
$P$ $=$ $T_{2}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{2},v_{2})}$
$\vdots$
$P$ $=$ $T_{n}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{n},v_{n})}$
where $\{T_{i}^{camera\leftarrow tool}|i=1,...,n\}$ denote the corresponding transformation matrix for each 2D image location of the pinhead, from tracker tool coordinate system to camera coordinate system, recorded from optical camera. The complete calibration matrix is expressed as $T_{calib}=T_{rotation}\cdot T_{scale}$. This composition ensures that pixels are first scaled into physical space, and then mapped into the tracker tool coordinate system using a rigid transformation.
Specifically, in each ultrasound image plane, the 2D location of the pinhead $p_{(u_{i},v_{i})}$ is manually identified in the image coordinate system, while its corresponding physical position in the camera coordinate system, denoted as $P$, remains constant but unknown throughout the acquisition. Consequently, in Eq. (3.2), the parameters to be estimated include the scaling factors ($s_{x}$ and $s_{y}$), the 6-DoF (three rotation angles and three translation components) comprising the rigid transformation $T_{rotation}$, and the 3D position of $P$. These parameters are jointly estimated using a nonlinear least-squares optimisation algorithm [35], in which the objective is to minimise the distance between the transformed 3D locations and the estimated fixed 3D location of the pinhead. The optimisation formulation is given as:
$\mathcal{L}$ $=$ $\min_{T_{rotation},\,T_{scale},\,P}\,\sum_{i=1}^{n}\,dist\,(T_{i}^{camera\leftarrow tool}\cdot T_{rotation}\cdot T_{scale}\cdot p_{(u_{i},v_{i})},P)$ (20)
$=$ $\min_{\mathbf{R},\,\mathbf{t},\,s_{x},\,s_{y},\,x,y,z}\,\sum_{i=1}^{n}\,dist\,(T_{i}^{camera\leftarrow tool}\cdot\left[\begin{array}[]{cc}\mathbf{R_{3\times 3}}&\mathbf{t_{3\times 1}}\\ \mathbf{0}&1\end{array}\right]\cdot\left[\begin{array}[]{cccc}s_{x}&0&0&0\\ 0&s_{y}&0&0\\ 0&0&1&0\\ 0&0&0&1\end{array}\right]\cdot\left[\begin{array}[]{cccc}u_{i}\\ v_{i}\\ 0\\ 1\end{array}\right],\left[\begin{array}[]{cccc}x\\ y\\ z\\ 1\end{array}\right])$
where $dist\,(\cdot)$ denotes the Euclidean distance computed between corresponding pairs of transformed 3D coordinates and the fixed 3D location $P$.
3.3 Transformation Estimation
The learning-based freehand ultrasound reconstruction task can be formulated as a pose regression problem, where the goal is to estimate transformations directly from ultrasound frames. The network architecture can vary in terms of its input configuration (e.g., frame pairs or sequences), supervisory label / output representation (rigid or non-rigid; 6-DoF, $4\times 4$ transformation matrix, or points coordinates), and loss function design (in Euclidean space or parameter space).
Input configuration. Given an ultrasound scan, the network can process a sequence of ultrasound frames to estimate the relative transformations between them. A common approach is to use two adjacent frames as input, which can be seen as a specific case of the general sequence input, and predict the relative transformation between them. The 3D ultrasound volume can then be reconstructed by accumulating these estimated relative transformations, as described in Eq. (4). Alternatively, the network can estimate the transformation between non-adjacent frames. This approach enables the network to leverage long-term spatial dependencies, potentially improving robustness and reducing cumulative error in trajectory estimation. A limitation of this approach is that it may not provide transformation estimates for all frames in the sequence.
Supervisory label / Output representation. Since the transformation from the tracker tool to the camera coordinate system, $T_{i}^{camera\leftarrow tool}$, is defined relative to the camera pose, it depends on the external configuration of the camera. This means that scanning the same object from different camera poses produces different transformations (from the tracker tool to the camera coordinate system). As a result, the same input content can be associated with different supervisory labels, which introduces ambiguity and hinders model generalisation. Therefore, transformations that are invariant to the camera’s pose are generally preferred in this application. The supervisory label can be designed as the transformation between two tracker tool coordinate systems corresponding to two ultrasound frames, denoted as $T_{j\leftarrow i}^{tool}$. Alternatively, the calibration matrix can be incorporated into labels to express the transformation in physical space (e.g., in millimeters), $T_{j\leftarrow i}$, or pixel space, via the scaled transformation $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$. It is important to note that all the three supervisory labels, $T_{j\leftarrow i}^{tool}$, $T_{j\leftarrow i}$, and $T_{scale}^{-1}\cdot T_{j\leftarrow i}\cdot T_{scale}$ are rigid transformations. It is crucial to highlight the relationship between the supervisory labels and evaluation metrics discussed in Section 4.3.1, where the supervisory labels serve as ground truth for model training, and the evaluation metrics are employed to assess model performance.
Other examples of the supervisory labels are 6-DoF vector, consisting of three rotation and three translation parameters, and seven-parameter representation using quaternion, comprising four rotation parameters and three translations. These representations are typically derived from the $4\times 4$ rigid transformation matrices, as mentioned above. Additionally, point coordinates derived from Eq. (3) can also be used as supervisory labels. Specifically, the baseline method of this Challenge used the 6-DoF vectors for supervision. It is worth noting that the representation of the network output can be similar to, but does not need to match, the format of the labels used during training. For example, the network may be designed to output a 6-DoF vector, even if the ground truth labels are provided in the form of transformation matrices.
Loss functions. Examples of the loss function include the Euclidean distance between predicted and ground truth point coordinates, as well as the difference between parameters or transformation matrices, regardless of the specific types of labels or predictions. For example, if the ground truth is provided as a $4\times 4$ transformation matrix and the network predicts a 6-DoF vector, the loss can be computed based on the difference between the predicted and ground truth 6-DoF parameters, where the latter are derived from the transformation matrix. Alternatively, the loss can be defined as the point-wise Euclidean distance between points transformed by the ground truth matrix and those transformed by the prediction.
3.4 Three-dimensional Reconstruction
To reconstruct the entire scan into 3D space, the relative transformation between each frame and a reference frame must be known. This can be achieved either by directly estimating the relative transformations with respect to the reference frame, or by estimating the transformations between pairs of consecutive frames (or non-adjacent frame pairs) and accumulating them accordingly. The reconstruction can be considered complete once all frame positions are estimated in a common reference coordinate system. Although full 3D volume reconstruction, such as interpolating scattered pixel intensities onto a regular voxel grid using methods like nearest-neighbor, linear, or weighted interpolation [15], is useful in some applications, it is not essential for many clinical applications and falls outside the scope of this Challenge.
4 Challenge Design
The TUS-REC2024 Challenge [^1] https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/ [^2] https://doi.org/10.5281/zenodo.10991500 is designed following the BIAS [36] Reporting Guideline for enhanced quality and transparency of biomedical research. This Challenge is associated with 5th International Workshop of Advances in Simplifying Medical UltraSound (ASMUS) at MICCAI 2024. The training and validation datasets are publicly available under CC BY-NC-SA license. The Challenge is an open-ended Challenge, and submissions are welcome even after the official deadline. The test set remains held out and will be used exclusively for benchmarking reconstruction performance.
4.1 Task Description
Aiming at estimating the location for each ultrasound frame in 3D space, this Challenge is tasked to predict four different sets of transformation-representing dense displacement field (DDF), a set of displacement vectors on individual pixels and a set of displacement vectors on provided landmarks, at both global and local levels. The global level DDFs denote the displacement between the current frame and the first frame, and the local level DDFs represent the displacement between the current frame and the previous frame. There are no restrictions on the internal design of the algorithm, for example, whether it is learning-based; processes data at the frame, sequence, or scan level; or assumes rigid, affine, or non-rigid transformations.
Participating teams are provided with sequential data and may choose to leverage its spatiotemporal information if useful. Each team’s model should take an ultrasound scan as input and output four sets of pixel-wise displacement vectors, representing the transformations to a reference frame (i.e., the first frame or the previous frame in the sequence). During evaluation, the submitted dockerized models will be used to generate these displacement fields, from which accuracy scores will be computed to assess reconstruction performance at both local and global levels.
4.2 Dataset
4.2.1 Data Collection
The dataset [^3] https://doi.org/10.5281/zenodo.11178508444https://doi.org/10.5281/zenodo.11180794555https://doi.org/10.5281/zenodo.11355499666https://doi.org/10.5281/zenodo.12752245 used in this Challenge was collected from both the left and right forearms of 85 volunteers at University College London (UCL), United Kingdom. This study was performed in accordance with the ethical standards in the 1964 Declaration of Helsinki and its later amendments or comparable ethical standards. Approval was granted by the Ethics Committee of local institution on 20th Jan. 2023 [24055/001]. The subject cohort was diverse in terms of race, gender, and age. Fig. 2 illustrates the equipment setup used during data acquisition. There were no specific exclusion criteria, except for individuals with allergies or skin conditions that could be aggravated by the ultrasound gel. All scanned forearms were confirmed to be in healthy condition.
Figure 2: Experimental setup for freehand ultrasound data acquisition. The setup consists of a tracked ultrasound probe, an ultrasound scanner, an optical tracker, and an acquisition laptop. The optical tracker monitors the probe’s transformation during scanning, while a volunteer is scanned using predefined probe trajectories.
2D ultrasound images were acquired using an Ultrasonix machine (BK, Europe) equipped with a curvilinear probe (4DC7-3/40). The ultrasound frames were captured at a rate of $20$ frames per second, with a resolution of $480\times 640$ pixels, without speckle reduction. Imaging was performed at a frequency of $6$ MHz, with a dynamic range of $83$ dB, an overall gain of $48\%$, and a depth of $9$ cm. Both the left and right forearms of volunteers were scanned. For each forearm, the ultrasound probe was moved along three trajectories, Straight line shape, C shape, and S shape, in both distal-to-proximal and proximal-to-distal directions. These scans were performed with the ultrasound plane perpendicular of and parallel to the scanning direction.
The selection of an appropriate tracking system is determined by the specific clinical environment and the required level of spatial accuracy. In this study, an optical tracking system was chosen due to its greater accuracy and operational convenience compared to EM tracking systems. Specifically, the NDI Polaris Vicra (Northern Digital Inc., Canada) was employed. During acquisition, position data recorded by the optical tracker was captured by the PLUS Toolkit [30], alongside the ultrasound images. For each ultrasound frame, the system records the pose of the tracker tool in the camera coordinate system, represented as a homogeneous transformation matrix. The calibration matrix was obtained using a pinhead-based method [37], defining the transformation between the ultrasound image coordinate system and the tracker tool coordinate system at the time of data acquisition. The data is temporally calibrated, with timestamps aligned between the transformations recorded by the optical tracker and the ultrasound frames acquired from the ultrasound machine, as described previously in Section 3.2.
4.2.2 Sources of Errors
The primary source of error arises from the precision limitations of the optical tracker. All labels were obtained using an optical tracker with the 3D root-mean-square (RMS) volumetric accuracy acceptance criterion being less than or equal to 0.25 mm and the 3D RMS repeatability acceptance criterion being less than or equal to 0.20 mm. Slight forearm movements may occur during scanning, which is also expected in clinical environments where this technique would be deployed. These motion-induced errors are assumed to be random across different cases. Their impact will be accounted for in the statistical analysis during result summarisation and interpretation. Additional sources of error include inaccuracies in the calibration process (both spatial and temporal), pressure-induced skin deformation, as well as the intra- and inter-observer variability during ultrasound data acquisition, affecting probe positioning and image capture quality. These levels of error are much lower than the typical state-of-the-art reconstruction errors in this application, which has been widely reported to range from several millimeters to tens of millimeters.
4.2.3 Data Pre-processing
For each scan, ultrasound frames with invalid transformation matrices, which were typically caused by blocked line of sight, were excluded. The remaining raw images, along with their corresponding transformation matrices, were temporally ordered and stored as key-value records in a .h5 file.
4.2.4 Data Split
Statistical power analysis was performed to determine the appropriate test sample size and minimise the likelihood of Type I and Type II errors in hypothesis testing. The effect size was calculated using Cohen’s D value, where the system error of the optical tracker ($0.25$ mm) was considered the meaningful difference between group means, and the standard deviation ($0.46$ mm) was derived from the results reported by [38]. A statistical power analysis for a t-test, assuming a significance level of $0.05$ and a statistical power of $0.9$, indicated a required test sample size of $31$. To ensure adequate power, we rounded up to $32$ samples ($768$ scans in total). This setup limits the probability of a Type II error to $10\%$ and a Type I error to $5\%$.
The dataset was randomly split into training, validation, and test sets, comprising $50$, $3$, and $32$ subjects, respectively. This corresponds to $1200$, $72$, and $768$ scans, totaling 606597, 34746, and 384105 frames. Ultrasound scans from the same subject will be assigned to the same set which avoids the information leak. Detailed information is described in Table 2. Specifically, the structure of the validation dataset is the same as that of the test set to ensure compatibility with the pre-defined folder hierarchy and naming conventions. This design allows submitted Docker images to run seamlessly on the test set and also supports their use for parameter tuning during validation.
No specific constraints are imposed on the use of the training and validation datasets. For example, participants are free to use all data from both sets for model training, or they may split the training set into training, validation, and test subsets for parameter tuning. Additionally, the use of both public and private data is permitted, but participants must disclose any external data sources they utilise.
Table 2: Overview of the freehand ultrasound dataset used in the TUS-REC2024 Challenge. The table summarises the number of subjects, scans, and frames across the training, validation, and test sets, categorised by scan trajectory shapes (Straight line shape, C shape, S shape), scanning directions (Parallel vs. Perpendicular, Distal-to-proximal vs. Proximal-to-distal), and scanned arms (Left arm vs. Right arm).
Train Validation Test
Subjects 50 3 32
Scans 1200 72 768
Frames 606597 34746 384105
Straight line shape subjects 50 3 32
C shape subjects 50 3 32
S shape subjects 50 3 32
Straight line shape scans 400 24 256
C shape scans 400 24 256
S shape scans 400 24 256
Straight line shape frames 192117 10515 119421
C shape frames 202654 11655 128721
S shape frames 211826 12576 135963
Parallel scanning subjects 50 3 32
Perpendicular scanning subjects 50 3 32
Parallel scanning scans 600 36 384
Perpendicular scanning scans 600 36 384
Parallel scanning frames 298722 17228 188399
Perpendicular scanning frames 307875 17518 195706
Left arm subjects 50 3 32
Right arm subjects 50 3 32
Left arm scans 600 36 384
Right arm scans 600 36 384
Left arm frames 301155 17081 192118
Right arm frames 305442 17665 191987
Distal-to-proximal scanning subjects 50 3 32
Proximal-to-distal scanning subjects 50 3 32
Distal-to-proximal scanning scans 600 36 384
Proximal-to-distal scanning scans 600 36 384
Distal-to-proximal scanning frames 298803 16908 181844
Proximal-to-distal scanning frames 307794 17838 202261
4.3 Evaluation Metrics
4.3.1 Metrics Definition
We use DDFs to evaluate the reconstruction performance, borrowing the widely recognised term used in non-rigid image registration for clarity and intuition. For each scan, participating methods are tasked to generate two types of displacement vectors representing frame-to-frame transformations, hereinafter referred to as predictions, at both global and local levels: 1) global displacement vectors are used to reconstruct all frames (excluding the first) relative to the first frame of the scan, which serves as the global reference frame; 2) local displacement vectors are used to reconstruct each frame (excluding the first) relative to its immediately previous frame, which serves as the local reference frame.
The performance of each submitted method will be assessed for every scan using two metrics: landmark reconstruction error and pixel reconstruction error: 1) landmark reconstruction error is defined as the average Euclidean distance between the ground-truth-reconstructed frame and the prediction-reconstructed frame, computed over a predefined set of landmarks. 2) pixel reconstruction error is similarly defined as the average Euclidean distance between the ground-truth and predicted reconstructions, calculated over all pixels in every frame except the first.
Accordingly, each method should produce the following four sets of displacement vectors:
  • Global-Pixel (GP) vectors – one per pixel (excluding the first frame) for global-level pixel reconstruction;
  • Global-Landmark (GL) vectors – one per landmark for global-level landmark reconstruction;
  • Local-Pixel (LP) vectors – one per pixel (excluding the first frame) for local-level pixel reconstruction;
  • Local-Landmark (LL) vectors – one per landmark for local-level landmark reconstruction.
Based on these outputs, four evaluation metrics will be computed:
  • Global Pixel Reconstruction Error (GPE) – the pixel reconstruction error calculated using GP vectors;
  • Global Landmark Reconstruction Error (GLE) – the landmark reconstruction error calculated using GL vectors;
  • Local Pixel Reconstruction Error (LPE) – the pixel reconstruction error calculated using LP vectors;
  • Local Landmark Reconstruction Error (LLE) – the landmark reconstruction error calculated using LL vectors.
Specifically, runtime will be included as an additional evaluation metric. It is defined as the consumed time of predicting the positions for all frames but the first frame in a scan, averaged across all scans in the test set. The scale-invariant feature transform (SIFT) [39] algorithm was applied to detect landmarks. For each scan, 20 landmarks with the highest response values were selected.
4.3.2 Rationale of Evaluation Metrics
Use of Euclidean distance-based error metrics vs. transformation parameter-based errors. Direct evaluating the accuracy of parameters of transformation matrix presents challenges, as the relative influence and weighting of rotational and translational components can vary significantly depending on experimental setups, imaging configurations, reference coordinate systems, and definitions of rotational axes. These factors are also often application-dependent. Therefore, this Challenge adopts Euclidean distance-based metrics, which offer a more direct and practical assessment of the discrepancy between ground truth and predicted positions in physical space.
Use of displacement-based transformation representations vs. rigid / affine matrices. Although ground-truth transformations are provided in the form of rigid transformation, we argue, based on practical experience in developing similar numerical algorithms, that requiring submissions to output homogeneous transformation matrices is not only unnecessary, but sometimes misleadingly encourages a more numerically challenging solution due to issues such as gimbal lock in using rotation matrix, local minima in numerical optimisation. In contrast, displacement-based representations allows flexibility for a quantitatively more accurate reconstruction, with a near-rigid transformation, which may be clinically sufficient [15]. Importantly, there are no restrictions on the internal methodology: participants may choose to internally estimate a rigid transformation matrix and convert it into the four required displacement vector sets for submission.
Justification for local and global reconstruction error metrics. Local and global reconstruction errors capture complementary aspects of algorithm performance. Global reconstruction (relative to the first frame) can reveal accumulated drift over time, while local reconstruction (relative to the immediately previous frame) assesses frame-level reconstruction. These metrics are therefore indicative of both short- and long-term accuracy. Although other monotonic metrics such as final drift and Dice overlap are commonly used [7], they are excluded here to streamline evaluation. In practice, one might choose to reconstruct a sequence of ultrasound frames (as opposed to the entire scan or two adjacent frames, which are represented by local and global errors, respectively), using a pre-optimised sequence length that is most suitable to the downstream application. Since this Challenge is designed without targeting a specific clinical use case, both local and global reconstruction errors are included to span the spectrum of reconstruction performance and provide a comprehensive assessment of algorithmic accuracy.
4.4 Ranking Scheme
The ranking follows the “aggregate then rank” strategy [40]. For each test scan, the four reconstruction error metrics will be normalised to the range $[0,1]$ using the formulas below.
$GPE^{*}$ $=(GPE_{max}-GPE)/(GPE_{max}-GPE_{min})$
$GLE^{*}$ $=(GLE_{max}-GLE)/(GLE_{max}-GLE_{min})$
$LPE^{*}$ $=(LPE_{max}-LPE)/(LPE_{max}-LPE_{min})$
$LLE^{*}$ $=(LLE_{max}-LLE)/(LLE_{max}-LLE_{min})$ (21)
where the superscript denotes the normalised reconstruction error, and the subscript min and max denote the minimum and maximum errors among all participats submissions, for each corresponding metric. For each scan, the final score is computed as a weighted average of the four normalised metrics:
$final\,score=0.25\times GPE^{*}+0.25\times GLE^{*}+0.25\times LPE^{*}+0.25\times LLE^{*}$ (22)
Each team’s overall score was calculated as the average final score across all test scans. This score, ranging from 0 to 1, determines the final ranking of all submitted algorithms. Scores were reported to three decimal places, with higher values indicating better performance.
For further insight, we also reported four other categories of scores, for reference and research interest without formal ranking: global reconstruction score = $0.5\times GPE^{*}+0.5\times GLE^{*}$, local reconstruction score = $0.5\times LPE^{*}+0.5\times LLE^{*}$, landmark reconstruction score = $0.5\times GLE^{*}+0.5\times LLE^{*}$ and pixel reconstruction score = $0.5\times GPE^{*}+0.5\times LPE^{*}$.
All evaluation metrics are normalised to a common scale to prevent metrics with inherently larger magnitudes from disproportionately influencing the overall score. The two levels of measurement (global and local) and the two types of displacement vectors (pixel-based and landmark-based) are considered equally important in achieving desirable reconstruction performance. Accordingly, equal weighting is applied to each metric to establish a fair and balanced benchmark for the Challenge. A minimum score of 0 was assigned to any case where the submitted code failed to execute or the evaluation metrics cannot be computed successfully. In the event of tied overall scores, ranking was determined based on runtime. A smaller runtime was awarded a higher rank. To encourage usability in the clinical applications, a maximum runtime limit of 2 minutes per scan was enforced for all Challenge submissions. Additionally, the raw (unnormalised) values of all defined evaluation metrics were made publicly available for transparency and further analysis.
4.5 Validation and Submission
A small validation set was provided to allow participants to tune their models on previously unseen data. An example Docker template [^7] https://github.com/QiLi111/tus-rec-challenge_baseline/tree/main/submission for evaluation on the validation dataset was provided, along with the corresponding evaluation metrics. This serves to enhance the validity of the submitted Docker images and improve overall transparency in the evaluation process. Docker images from participating teams were submitted via an online form, which included a brief method description and step-by-step instructions for downloading and executing the Docker image. All submitted methods must operate in a fully automatic manner. Participants are permitted to modify and resubmit their Docker image if it fails to run on the test set due to issues such as incorrect input / output formatting or data format incompatibility. Each team was allowed a maximum of five submissions, provided the submissions represent substantively different approaches rather than minor variations in hyperparameters. The best-performing result among these were considered as the team’s final result. All submitted Docker images were independently tested by two members of the Challenge organisition team using the hidden test dataset. Evaluations were conducted on two separate platforms with identical hardware configurations: Ubuntu 18.04.6 LTS, Intel(R) Xeon(R) Gold 5215 CPU @ 2.50GHz (20 cores), NVIDIA Quadro GV100 GPU (32GB VRAM), and 128GB RAM.
4.6 Awards
Results from all participants were publicly displayed on the official leaderboards, except in cases where submissions encountered errors during the evaluation process. Additional certificates of recognition were awarded to the first-place team and the runner-up. All teams with successfully evaluated submissions received certificates of participation.
4.7 Timeline
The TUS-REC2024 Challenge is an open-call event designed to encourage broad community participation and, although this edition was structured as a one-time event tied to MICCAI 2024, its infrastructure and open-submission framework support potential future iterations, enabling continued engagement beyond the initial evaluation cycle.
The official timeline is aligned with MICCAI 2024, as detailed in Fig. 3. The Challenge began with the website launch and team registration on April 1, 2024, followed by the release of training data on May 13, and baseline code on June 23. On July 29, validation data and the Docker image template were released, offering participants a clear structure for submission. This template, along with an evaluation script that incorporates the Challenge metrics, aimed to ensure transparency and reproducibility in assessment and was designed to align with the BIAS Reporting Guideline. The submission window officially opened on August 12 and closed on September 9. The announcement of the winning teams took place on September 16, and the TUS-REC2024 Challenge event was held on October 6, 2024, during MICCAI 2024.
Figure 3: Timeline of the TUS-REC2024 Challenge. Key milestones include the release of training data, baseline code, and validation resources, followed by the submission phase and final Challenge event at MICCAI 2024.
Figure 4: Participant and team statistics summarising engagement across TUS-REC2024 Challenge, including 101 registered participants from 43 teams, and participation by 25 individuals grouped into 6 teams.
5 Challenge Outcome
5.1 Participation Statistics
Fig. 4 presents the participant statistics for TUS-REC2024 Challenge. By the submission deadline, a total of 101 individuals registered, representing 43 teams comprising members from both academia and industry. Participants came from 14 countries across 5 continents, reflecting the international interest and global reach of the Challenge. Despite strong initial engagement, six teams submitted their Docker images by the submission deadline, involving a total of 25 participants. In total, 21 valid docker images were received. The number of submissions varied across teams, with several teams submitting multiple Docker images for performance optimisation. This decline from registration to submission may reflect the technical complexity of the task or limited preparation time. Notably, the majority of registered and submitting teams were affiliated with academic institutions, particularly universities. Overall, the statistics highlight both the broad appeal of the Challenge, and the practical hurdles faced by participants in progressing from initial registration to successful submission, such as time constrains and difficulties in model development.
5.2 Methods
5.2.1 Methodologies of Baseline and Participating Teams [^8] This is a summary of TUS-REC2024 Challenge, rather than proposing these methods. The authors may publish their own technical papers enabling reproducibility of their methods.
This section presents the approaches of the top five participating teams, alongside the baseline approach provided by the organisers. Table 3 summarises key information about these teams, including their model abbreviations, methodological highlights, team names, and institutional affiliations. Table LABEL:imp summarises the implementation details of the baseline and top five participating methods. It includes model architectures, backbones, training configurations, loss functions, and other relevant technical aspects that highlight the diversity of approaches adopted in the Challenge.
Table 3: Overview of the top five participating teams in TUS-REC2024 Challenge, including model abbreviations, methodological descriptions, team names, and institutional affiliations.
Rank Model Abbreviation Method Team Name Affiliation(s)
1 FiMoNet Enhanced Fine-grained Motion Network MUSIC Lab Shenzhen University; Shenzhen RayShape Medical Technology Inc.
2 RecuVol Recurrent CNN-LSTM Trackerless Freehand 3D Ultrasound Reconstruction ISRU@DKFZ DKFZ (German Cancer Research Center) Heidelberg; University of Cincinnati; Tufts University
3 FlowNet Three-dimensional Ultrasound Reconstruction using CNN Learned by Flow Field Transformation zjr Hong Kong Centre for Cerebro-cardiovascular Health Engineering; City University of Hong Kong
4 MoGLo-Net Motion-based Learning Networks with Global-Local Attention for Ultrasound Scan Motion Estimation AMI-Lab Pusan National University
5 PLPPI Physics Guided Learning-based Prediction of Pose Information UW-Madison Elastography Lab University of Wisconsin-Madison
5.2.1.1 Baseline Algorithm [^10] https://github.com/QiLi111/tus-rec-challenge_baseline
The baseline method utilises the EfficientNet-B1 architecture [41], taking as input a pair of adjacent ultrasound frames. The network predicts a 6-DoF transformation, representing the transformation from the image coordinate system (in mm) of one frame to that of the other. The training loss is formulated as the mean squared error (MSE) between point coordinates transformed by the ground truth and those transformed by the network’s prediction.
$\mathcal{L}=D\,(T^{gt}_{j\leftarrow i}\cdot T_{scale}\cdot\mathbf{p}_{corner},\,T_{j\leftarrow i}\cdot T_{scale}\cdot\mathbf{p}_{corner}),\,\,i=j+1$ (23)
where $D(,\cdot,)$ denotes the MSE computed over the $x$, $y$ and $z$ coordinates of corresponding points. $T^{gt}_{j\leftarrow i}$ and $T_{j\leftarrow i}$ represent the ground truth and predicted transformation matrices, obtained by converting the predicted 6-DoF into a homogeneous transformation matrix. $T_{j\leftarrow i}$ is defined as in Eq. (1). $\mathbf{p}_{corner}$ denotes the coordinates of the four corner pixels in the image coordinate system (in pixels).
During inference, the 6-DoF between adjacent frames are estimated by sequentially inputting frame pairs into the network. The local DDFs for all pixels within an image are computed by left-multiplying the predicted local transformation matrix with the calibrated image coordinates and subtracting the coordinates of the reference frame: $DDF_{local}^{(i)}=T_{local}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$, where $\mathbf{p}$ denotes the coordinates of all pixels within an image in the image coordinate system (in pixels). $T_{local}^{(i)}=T_{i-1\leftarrow i}$ is the transformation matrix from frame $i$ to frame $i-1$, converted from the 6-DoF. The global transformation from any frame $i$ to the first frame, $T_{global}^{(i)}$, is derived by composing predicted local transformations through left-multiplication in reverse temporal order: $T_{global}^{(i)}=T_{local}^{(2)}\cdot T_{local}^{(3)}\cdots T_{local}^{(i)}$. The global DDFs for all pixels within an image are computed using $DDF_{global}^{(i)}=T_{global}^{(i)}\cdot T_{scale}\cdot\mathbf{p}-T_{scale}\cdot\mathbf{p}$. The local and global DDFs at predefined landmark locations can be obtained either by indexing the corresponding positions from $DDF_{local}$ and $DDF_{global}$, respectively, or by calculating them using the formula above, replacing all pixel coordinates $\mathbf{p}$ with the landmark locations.
5.2.1.2 FiMoNet
Fine-grained spatio-temporal learning is essential for freehand 3D ultrasound reconstruction. To address the complexities of long-range dependencies introduced by diverse probe motions as well as the large number of patches involved in spatio-temporal modeling, we adapted Mamba [42]. Mamba utilises the state-space model’s capacity to manage long-range dependencies, providing an effective solution for this task.
We employed ensemble learning to combine two models:
  • Model 1 consists of ResNet18 and ReMamba [43]. Following the method in [43], convolutional blocks from ResNet18 and ReMamba blocks are applied to extract fine-grained image features at multiple scales. A fully connected layer is then employed to regress the 6-DoF transformation parameters.
  • Model 2 integrates ResNet18 with a multi-layer Mamba block. Inspired by [24], a cascaded architecture is designed. Specifically, the final fully connected layer of ResNet18 is removed and replaced with a multi-layer Mamba block, followed by a fully connected layer to produce the output.
For the 6-DoF transformations between adjacent frames, estimated by the network as $\theta$ and the ground truth 6-DoF transformation $\theta^{gt}$, we employed both the $L_{1}$ loss and Pearson correlation loss:
$\mathcal{L}=\parallel{\theta^{gt}}-\theta\parallel_{1}+\left(1-\frac{Cov(\theta^{gt},\theta)}{\sigma(\theta^{gt})\sigma(\theta)}\right)$ (24)
where $Cov(\theta^{gt},\theta)$ represents the covariance between ground truth and predicted 6-DoF parameters. $\sigma(\cdot)$ denotes the standard deviation. During inference, the model takes the entire scan as input and outputs the 6-DoF transformation between all adjacent frames. These local transformations are then converted to global 6-DoF transformations, which are used to generate the global DDFs.
5.2.1.3 RecuVol [^12] https://github.com/ISRU-DKFZ/RecuVol
The proposed approach utilises an EfficientNet-based CNN (pre-trained on ImageNet) to extract features from pairs of consecutive frames. These features are processed sequentially by a LSTM network to model temporal dependencies. The network predicts 3D translation and rotation parameters for each frame pair. Training is performed by minimising the MSE loss on these parameters, enabling the model to learn robust frame-to-frame alignments. TrivialAugment is used for data augmentation, and sequences of 16 frames are processed at a time, with adjacent frames concatenated prior to input into the CNN.
We employed a 5-fold cross-validation strategy for training; however, one fold displayed instability and was consequently excluded. The remaining four folds were ensembled by computing the median of the predicted 6-DoF transformation parameters, yielding a single final prediction. To further enhance the performance, a second 4-fold ensemble was trained on data downsampled by a factor of 1.25. The final submission is composed of both 4-fold ensembles (original and downsampled), resulting in a total of eight models.
During inference, the model estimates the rigid transformation parameters between each consecutive pair of frames within a scan. By sequentially concatenating these pairwise transformations starting from the first frame, we compute the global pose of each frame relative to the first frame. Using all frames’ global transformations, the 3D volume is reconstructed. The DDF is derived by back-mapping voxel coordinates from a reference 3D grid to their original frame positions. The model only explicitly predicts local transformations, while global transformations are obtained by sequentially accumulating these local estimates. Both local and global transformations are rigid and derived from the model’s frame-to-frame predictions.
5.2.1.4 FlowNet
The network is based on EfficientNet-B6, taking $n=10$ consecutive ultrasound frames $S\in\mathbb{R}^{n\times h\times w}$ as input. It outputs a set of transformation parameters $Y\in\mathbb{R}^{(n-1)\times 6}$, where each 6-DoF vector represents the rigid transformation from the last ultrasound frame $S^{(n)}$ to a preceding frame $S^{(i)},\,i\in[1,n-1]$. $Y$ is used to compute a flow field $F$, enabling the warping of $n-1$ frames to generate $S^{warp}\in\mathbb{R}^{(n-1)\times h\times w}$. $Y$ can also be expressed as matrices $T_{Y}$, where each $T_{Y}^{(i)}$ denotes the transformation matrix from $S^{(n)}$ to $S^{(i)}$. The transformation matrix between any two frames $S^{(i)}$ and $S^{(j)}$ can be obtained by calculating $T_{Y}^{(j)}\cdot(T_{Y}^{(i)})^{-1}$, forming the dense transformation matrix set $T^{ds}$. The resulting dense point coordinates $P^{ds}$ are then used to calculate the overall loss:
$\mathcal{L}=\text{MSE}\left(P_{gt}^{ds},P^{ds}\right)+0.5\times\text{MSE}\left(T_{gt}^{ds},T^{ds}\right)+0.5\times\text{MSE}\left(S,S^{warp}\right)$ (25)
where $P_{gt}^{ds}$ and $T_{gt}^{ds}$ denote the ground truth points coordinates and transformations, respectively.
Three models were selected: the final epoch model, the model from 100 epochs earlier, and the one with the lowest validation distance. Given the full scan $\mathcal{S}\in\mathbb{R}^{N\times h\times w}$, with local and global transformations $T_{local},T_{global}\in\mathbb{R}^{(N-1)\times 4\times 4}$, we sequentially process sequences of $n$ frames using a stride of $n-1$, such that the last frame of one sequence is the first of the next. Predictions from the three models are averaged to obtain the final $Y$ and $T_{Y}$. For the first sequence, local transformation could be calculated by $T_{local}^{(i)}=T_{Y}^{(i-1)}\cdot(T_{Y}^{(i)})^{-1}$, and global transformation is calculated by $T_{global}^{(i)}=T_{Y}^{(1)}\cdot(T_{Y}^{(i)})^{-1}$. After computing local transformations for all frames in the first sequence, subsequent sequences are processed sequentially using the same method to obtain local transformations. The global transformation for the $k^{th}$ frame in $l^{th}$ sequence $S^{((l-1)\times(n-1)+k)}$ is computed as $T_{global}^{((l-1)\times(n-1)+k)}=T^{(1)}_{Y,s_{1}}\cdot T^{(1)}_{Y,s_{2}}\cdots T^{(1)}_{Y,s_{l}}\cdot(T_{Y,s_{l}}^{(k)})^{-1}$, where $T^{(k)}_{Y,s_{l}}$ denotes the transformation matrix from $n^{th}$ frame to the $k^{th}$ frame in the $l^{th}$ sequence. This yields $T_{local}$ and $T_{global}$ for the full scan. The scan is then reversed, and the same procedure is applied to obtain $T_{local}^{reverse}$ and $T_{global}^{reverse}$. Final predictions are obtained by averaging the forward and reversed results: $T^{avg}_{local}=(T_{local}+T_{local}^{reverse})/2$, $T^{avg}_{global}=(T_{global}+T_{global}^{reverse})/2$.
To further improve local transformation prediction, the scan is offset by excluding the first $m=1,2,3,4$ frames, yielding sub-scans $\mathcal{S}_{m}\in\mathbb{R}^{(N-m)\times h\times w}$. Corresponding local transformations $T^{avg}_{local,m}$ are computed for each offset. The final local transformation is obtained by averaging all predictions: $T^{final}_{local}=(T^{avg}_{local}+T^{avg}_{local,1}+T^{avg}_{local,2}+T^{avg}_{local,3}+T^{avg}_{local,4})/5$.
5.2.1.5 MoGLo-Net [^14] https://github.com/guhong3648/US3D
Input images are cropped to square regions to remove background artifacts and normalised to $[-1,1]$. As shown in Fig. 5, we developed MoGLo-Net, a motion-based learning network with global-local attention. Two ultrasound sequences, each consisting of $k+1$ frames, are processed in parallel through a ResNet-based encoder for consistent feature refinement.
Figure 5: Overview of MoGlo-Net.
The correlation volumes $C$ are computed from encoded features of the two sequences. A patch-wise correlation operation, inspired by [1, 44, 45], models local relationships between successive frames by (1) defining a common region of interest (ROI) on adjacent feature maps, (2) computing cosine similarity between local patches by sliding one patch over the entire ROI, and (3) aggregating results from multiple ROIs to form the correlation volume. The correlation volume encodes motion cues, enhancing the model’s motion estimation accuracy.
The encoded features are merged into $M$ and refined with $C$ via encoder blocks (purple arrows). The resulting features are concatenated to form the final feature map $F$, where the global-local attention module is applied. This module is designed as a self-attention mechanism [46, 47, 48]: (1) global (GF) and local features (LF) are downsampled by factors of 4 and 2, respectively; (2) LFs are extracted from early encoder layers as patch-wise feature blocks; (3) both GFs and LFs are refined using a conventional attention mechanism; (4) cosine similarity between GFs and LFs serves as attention scores to weight LFs; and (5) weighted LFs are projected to aggregate local information. The final recalibrated feature, formed by concatenating GFs and LFs, is fed into the RNN-based estimator (Conv-GRU) to predict 6-DoF.
We employed three loss functions: Correlation Loss [11] ensures motion consistency; Triplet Loss [11] contrasts the final feature maps; and Motion-based Mean Absolute Error (MMAE) emphasises errors in fast-motion regions:
$\mathcal{L}_{\text{MMAE}}=\frac{1}{6(s+1)}\sum_{i=n}^{n+s}\sum_{k=1}^{6}w_{i}\left|\Delta{\theta}_{i,k}^{gt}-\Delta{\theta}_{i,k}\right|$ (26)
where $s+1$ denotes the number of frames in an ultrasound sequence. $w_{i}=\left|\Delta{\theta}_{i}\right|+\varepsilon$ is a motion-based weighting term. Fast motion errors are penalised more heavily, as $w_{i}$ increases with larger motion vectors. The smoothing term $\varepsilon$ prevents over-amplification. Our model predicts 6-DoF for ultrasound sequences, but only the final frame’s estimation is used during inference. Therefore, predicting the scan motion for the entire scan requires $N$ sequences, where $N$ is the number of frames in the scan. To do this, we add padding at the beginning of the scan frames. The global and local transformations are derived based on the TUS-REC2024 baseline code.
5.2.1.6 PLPPI [^16] https://github.com/Alphafrey946/PLPPI [^17] This work is summarised in [49], which provides further details on the method design.
To address the complexities in freehand ultrasound reconstruction, particularly out-of-plane motion [49], we proposed a lightweight, physics-informed deep learning model. Our dual-stream network decouples spatial and temporal learning, incorporating learnable operators to capture data priors for modeling temporal relationships and integrating a physical model to simplify learning, offering flexibility for various scanning paths.
The PLPPI model consists of spatial and temporal branches, followed by a fusion module and prediction head. The spatial branch uses 2D convolutions to aggregate intra-frame spatial context, while the temporal branch extracts inter-frame motion cues via speckle decorrelation patterns. This involves constructing a correlation volume to quantify the underlying motion information. The outputs from both branches are fused to represent the input sequence in the feature space, with speckle decorrelation serving as a key physics-based prior. The temporal branch computes a correlation volume $cv$ by measuring patch-wise similarity between two dense feature maps, $c_{21}$ and $c_{22}$, where $c\in\mathbb{R}^{h\times w\times d}$. $cv$ is defined as: $cv(x_{1},x_{2})=\sum_{s\in[-p,p]\times[-p,p]}c_{21}(x_{1}+s)^{T}\cdot c_{22}(x_{2}+s)$, with $x_{1}$, $x_{2}$ denoting the patch locations centered at $c_{21}$ and $c_{22}$, respectively, and $p$ the maximum displacement between $x_{1}$ and $x_{2}$. The squared patch size is $K=2p+1=21$. During training, input image stacks are split into two sub-volumes and passed through 2D convolutions to obtain $c_{21}$ and $c_{22}$. Temporal features are then bilinearly upsampled and fused with spatial features for joint representation.
Compared to our previous approach [49], we introduced two key modifications: (1) replacing the ResNet backbone with the pretrained foundation model Biomedical CLIP [50], and (2) redesigning the loss function to better leverage the capabilities of the foundation model. The new loss is defined as:
$\mathcal{L}=\alpha\parallel\theta^{gt}-\theta\parallel^{2}+\beta\parallel C({I^{gt}})-C(I^{recon})\parallel^{2}+\gamma\parallel{\theta^{gt}}\cdot p_{lmk}-\theta\cdot p_{lmk}\parallel^{2}$ (27)
The loss has three terms: (1) MSE between predicted and ground truth pose, (2) embedding consistency using Biomedical CLIP [50] on a “reconstructed” image $I^{recon}$, obtained by taking pixel-wise average of the two closest images to predicted ${\theta}$ and (3) projection loss as Euclidean distance between projected and true 3D landmarks, projected from 2D landmarks $p_{lmk}$. $\alpha$, $\beta$, $\gamma$ are hyperparameters. The Biomedical CLIP is finetuned with provided training data.
During inference, the model outputs $n-1$ local transformations estimated from $n$ input images. Sliding window averaging is applied to obtain the final local transfromations: ${\theta}_{local}^{(i)}=\frac{1}{W}\sum^{i}_{j=i-W+1}\theta^{(j)}$, with window size $W$. Global transformation is then computed as $T_{global}^{(i)}=\prod^{i}_{j=1}T_{local}^{(j)}$ where $T_{local}^{(j)}$ is converted from ${\theta}_{local}^{(j)}$.
5.2.2 Methodology Analysis Among Teams
Most of the proposed approaches leverage both spatial and temporal learning to capture long-term dependencies within ultrasound sequences. Examples include the use of Mamba modules in FiMoNet and LSTM networks in RecuVol. ResNet and EfficientNet are employed as backbone architectures across several methods. All models predict 6-DoF transformations. While most methods estimate frame-to-frame transformations, FlowNet predicts transforma-tions between non-adjacent (interval) frames. Regarding loss functions, the primary objective across methods is to
Table 4: Implementation details of the baseline and top five participating methods, including model architectures, training setups, loss functions, and data processing strategies.
Model abbreviation Baseline FiMoNet RecuVol FlowNet MoGLo-Net PLPPI
Architecture 2D CNN 2D CNN; State Space Model 2D CNN (extracts features) followed by LSTM 2D CNN 2D ResNet; Conv-GRU 2D CNN
Backbone EfficientNet-B1 ResNet18; Mamba ResNet EfficientNet-B6 ResNet ResNet-50 from [51]
Input sequence length 2 Not fixed, depends on scan length $\sim$16 10 5 6
Output Rigid; 6-DoF of adjacent frames Rigid; 6-DoF of adjacent frames 6-DoF 6-DoF 6-DoF Rigid; 6 DoF (utilising the representation in [52])
Model size (number of parameters) $\sim$6.5e6 1.8e7 $\sim$1e7 4.1e7 3.3e7 4.6e7
Model weights initialisation Random initialisation ResNet (ImageNet-1K initialisation); Mamba (random initialisation) ImageNet initialisation for CNN Kaiming normal distribution Random initialisation Kaiming normal distribution
Pretraining N/A N/A ImageNet pretrained CNN backbone N/A N/A Biomedical CLIP
Train/Val/Test splits 3:1:1 5:1:4 5 fold cross validation 3:1:1 45:5:3 8:1:1
Pre-processing N/A Resize image to 50% width and height Normalising, downsampling by 1.25 (for half of the final ensemble models) Normalising to $[0,1]$ Cropping; scaling to $[-1,1]$ Fine-tuned Biomedical CLIP on the training dataset
Data augmentation N/A Randomly sampling scans at different intervals; randomly flipping scans PyTorch TrivialAugment Flip the order of consecutive frames N/A Adding Gaussian noise, random cropping
Data sampling N/A Randomly sampling scans of different lengths, ranging from 60 to 180 Sequences of 16 consecutive frames of the same scan N/A Randomly sample ultrasound sequence with $k+1$ frames N/A
External data N/A N/A N/A N/A N/A N/A
Loss MSE loss on transformed points coordinates $L_{1}$ loss, Pearson correlation loss MSE loss on transformation parameters MSE loss MMAE loss, Correlation loss and Triplet loss MSE loss, Consistency loss, Projection loss ($\alpha=1$, $\beta=0.69$, $\gamma=0.67$)
Optimiser Adam Adam Adam Adam AdamW AdamW
Other details (e.g., any specific technique used) N/A Multi-directional state space model [43] Out of 5 folds, one was withheld due to unstable training N/A Motion-based MAE; correlation operation; global-local attention Self-attention and shift-invariance [53]; Bayesian search [54] for hyperparameter tuning
minimise the discrepancy between the predicted and ground truth transformation parameters, commonly using $L_{1}$, MSE, or Pearson correlation-based losses (e.g., FiMoNet, RecuVol). Additional loss formulations are also utilised: MoGLo-Net employs a triplet loss; FlowNet incorporates a point-based loss on transformed coordinates and an MSE loss between original and warped ultrasound images; and PLPPI integrates embedding consistency loss and projection loss on landmarks. Pre-training is adopted by three approaches: FiMoNet and RecuVol use ImageNet-pretrained weights, while PLPPI employs foundation model Biomedical CLIP. Ensemble learning is another common strategy. FiMoNet combines two distinct models; RecuVol aggregates eight models derived from two rounds of 5-fold cross-validation; and FlowNet selects three models from different training epochs.
Illustrated in Table LABEL:imp, all methods are trained in an end-to-end manner and utilise offline inference. Most teams employs the Adam optimizer, with two opting for AdamW. A uniform base learning rate of 1e-4 is used across all submissions, though learning rate scheduling varies, including approaches such as StepLR, ReduceLROnPlateau, and cosine annealing with warmup. Training epochs span from under 100 to 13,400. Batch size varies between 1 and 32. The teams utilise a variety of GPU configurations for model training, including single-GPU setups with NVIDIA Quadro GV100, RTX 3090, 4090, and A6000, as well as a dual-GPU setup with A40s. Training times range from 1.2 to 9.7 GPU days, depending on resources and setup. None of the teams report the use of external data during training. Standard preprocessing steps, including scaling, cropping, and normalisation, along with data augmentation techniques such as temporal sampling and flipping, are consistently applied.
5.3 Results Analysis
5.3.1 Overall Performance
Tables 5 and 6 present the performance of each team, assessed using four evaluation metrics along with their corresponding normalised scores. Figs. 6(a) and 7 provide a graphical representation of the same data. The abbreviations FS, GS, LS, PS, and LMS refer to the final score, global score, local score, pixel score, and landmark score, respectively. The evaluation results demonstrate that composite metrics effectively capture the strengths and limitations of participating methods across multiple spatial levels. The top-performing methods are characterised by strong local accuracy, low landmark error, and balanced global performance, indicating the importance of combining spatial-temporal modeling and ensemble learning.
FiMoNet leads in 3 out of 4 normalised scores, particularly in local scores (LS: $0.951\pm 0.074$) and frame-to-frame accuracy (LPE: $0.097\pm 0.014$, LLE: $0.084\pm 0.019$), reflecting the advantage of its use of Mamba for temporal modeling, Pearson correlation-based loss, and dual-model ensembling. Its relatively low runtime also highlights a favorable balance between accuracy and efficiency. Close behind, RecuVol shows more balanced performance across global and pixel-level metrics (e.g., PS: $0.835\pm 0.131$, GLE: $5.978\pm 3.719$), but slightly lower local precision than FiMoNet. This indicates that while its LSTM-based temporal modeling and extensive ensemble setup (eight models) improve robustness, it may not capture local spatial structures as effectively. FlowNet, although achieving the lowest global errors (GPE: $5.970\pm 3.523$, GLE: $5.167\pm 3.682$), ranks lower in local-related metrics. This suggests that its interval-based frame prediction strategy and point/image-based loss functions capture coarse alignment well, but are less suited for precise local alignment. Its high inference time also presents a practical limitation.
MoGLo-Net and PLPPI show lower performance across all scores, with notably low landmark ($0.551\pm 0.270$ and $0.322\pm 0.240$) and global scores ($0.548\pm 0.322$ and $0.272\pm 0.302$). This suggests that their strategies, such as triplet loss in MoGLo-Net and embedding/projection losses in PLPPI, may not compensate for the lack of strong temporal modeling or ensemble learning. The Baseline model shows comparatively lower performance across all metrics, particularly in local alignment (LS: $0.056\pm 0.106$) and landmark localisation (LLE: $0.118\pm 0.031$).
The final score, which normalises performance based on global and local transfromations, on all pixel and landmark level errors, ranks FiMoNet ($0.852$) and RecuVol ($0.817$) highest, indicating superior overall accuracy. These methods employ temporal modeling (Mamba and LSTM, respectively), diverse loss functions, and ensemble strategies, suggesting that integrating spatial-temporal features with strong supervision contributes to consistent performance across all spatial scales. Disaggregated metrics reveal further insights. The global score, based on GPE and GLE, highlights models that excel in aligning entire ultrasound scan. FlowNet, despite ranking third overall, achieves the best GPE ($5.970$ mm), reflecting strong global transformation learning. However, its local score is substantially lower ($0.622$), indicating that precise local alignment is not guaranteed by low global error alone. In contrast, FiMoNet achieves the highest local score (LS: 0.951), suggesting its feature extraction strategy, fine-grained feature extraction at multiple scales, is particularly effective at capturing anatomical detail. The pixel-wise score and landmark score aggregate global-local accuracy at pixel and landmark levels respectively. FiMoNet and RecuVol lead in both scores, reflecting their attention to both dense field alignment and landmark-specific accuracy.
Table 5: Performance of participating teams expressed as normalised scores based on evaluation metrics. An upward arrow ($\uparrow$) denotes that higher values indicate better performance, while a downward arrow ($\downarrow$) indicates that lower values correspond to better performance. Values highlighted in bold represent the best-performing results for each score.
Rank Model Abbreviation FS ($\uparrow$) GS ($\uparrow$) LS ($\uparrow$) PS ($\uparrow$) LMS ($\uparrow$) Run Time (s) ($\downarrow$)
1 FiMoNet $\mathbf{0.852\pm 0.130}$ $0.753\pm 0.230$ $\mathbf{0.951\pm 0.074}$ $\mathbf{0.875\pm 0.122}$ $\mathbf{0.829\pm 0.148}$ $9.213\pm 1.153$
2 RecuVol $0.817\pm 0.140$ $0.790\pm 0.205$ $0.844\pm 0.153$ $0.835\pm 0.131$ $0.799\pm 0.169$ $17.173\pm 1.800$
3 FlowNet $0.754\pm 0.145$ $\mathbf{0.886\pm 0.182}$ $0.622\pm 0.169$ $0.757\pm 0.135$ $0.751\pm 0.175$ $46.956\pm 5.617$
4 MoGLo-Net $0.573\pm 0.240$ $0.548\pm 0.322$ $0.598\pm 0.246$ $0.595\pm 0.233$ $0.551\pm 0.270$ $16.964\pm 2.015$
5 PLPPI $0.303\pm 0.215$ $0.272\pm 0.302$ $0.334\pm 0.200$ $0.285\pm 0.209$ $0.322\pm 0.240$ $15.112\pm 1.656$
6 Baseline $0.146\pm 0.159$ $0.236\pm 0.273$ $0.056\pm 0.106$ $0.125\pm 0.148$ $0.167\pm 0.186$ $\mathbf{8.135\pm 0.996}$
Table 6: Performance of participating teams measured by evaluation metrics. An upward arrow ($\uparrow$) denotes that higher values indicate better performance, while a downward arrow ($\downarrow$) indicates that lower values correspond to better performance. Values highlighted in bold represent the best-performing results for each metric.
Rank Model Abbreviation GPE (mm) ($\downarrow$) GLE (mm) ($\downarrow$) LPE (mm) ($\downarrow$) LLE (mm) ($\downarrow$)
1 FiMoNet $7.191\pm 3.687$ $6.281\pm 3.812$ $\mathbf{0.097\pm 0.014}$ $\mathbf{0.084\pm 0.019}$
2 RecuVol $6.858\pm 3.526$ $5.978\pm 3.719$ $0.101\pm 0.016$ $0.088\pm 0.021$
3 FlowNet $\mathbf{5.970\pm 3.523}$ $\mathbf{5.167\pm 3.682}$ $0.111\pm 0.016$ $0.096\pm 0.022$
4 MoGLo-Net $9.388\pm 5.358$ $8.459\pm 5.699$ $0.112\pm 0.024$ $0.100\pm 0.033$
5 PLPPI $12.093\pm 4.460$ $10.366\pm 5.006$ $0.122\pm 0.019$ $0.107\pm 0.025$
6 Baseline $12.490\pm 5.462$ $11.129\pm 5.838$ $0.135\pm 0.024$ $0.118\pm 0.031$
The statistical testing results below demonstrate the validity and effectiveness of the evaluation metrics:
  • For the five normalised scores, in scan level, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between methods PLPPI and baseline ($p$-value $=0.035$) in global score, methods FlowNet and MoGLo-Net ($p$-value $=0.033$) in local score. In subject level, for the five normalised scores, all pairwise team comparisons yield $p$-values below 0.001, except for the comparison between methods FiMoNet and RecuVol ($p$-value $=0.003$) in final score, FiMoNet and RecuVol ($p$-value $=0.021$) in global score, PLPPI and baseline ($p$-value $=0.358$) in global score, FlowNet and MoGLo-Net ($p$-value $=0.087$) in local score, FiMoNet and RecuVol ($p$-value $=0.021$) in landmark score.
  • For the four error metrics, in scan level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between PLPPI and the baseline method ($p$-value $=0.037$) in GPE metric, and FlowNet and MoGLo-Net ($p$-value $=0.011$) in LPE metric. In subject level, all pairwise team comparisons result in $p$-values less than 0.001, except for the comparison between methods FiMoNet and RecuVol ($p$-value $=0.008$) in GPE metric, methods PLPPI and baseline ($p$-value $=0.348$) in GPE metric, methods FiMoNet and RecuVol ($p$-value $=0.021$) in GLE metric, methods PLPPI and baseline ($p$-value $=0.081$) in GLE metric, and methods FlowNet and MoGLo-Net ($p$-value $=0.043$) in LPE metric.
Figure 6: (a) Box plots illustrating the distribution of performance scores across all test cases for each evaluated method (FiMoNet, RecuVol, FlowNet, MoGLO-Net, PLPPI, and Baseline), expressed in four normalised scores (FS, GS, LS, PS, LMS). The central line in each box represents the median score, with box edges indicating the first and third quartiles, and whiskers extending to 1.5 times the interquartile range (IQR). (b) Pairwise statistical comparison of final scores across methods in subject level, expressed as $p$-values from significance tests. Values close to zero indicate statistically significant differences in performance.
Figure 7: (a): Box plots of GPE and GLE for each team. (b) Box plots of LPE and LLE for each team. In both subfigures, lower error values indicate better performance. The central line in each box represents the median value, with box edges indicating the first and third quartiles, and whiskers extending to 1.5 times the interquartile range (IQR).
Figure 8: (a) Ranking stability of participating teams using bootstrap sampling (2,000 bootstrap samples, each equal to test set size). The area of each bubble corresponds to the relative frequency of the respective ranks observed across bootstrap samples. The median rank for each algorithm is represented by a black cross. (b) Sampling distributions of the mean final score for each algorithm, approximated using the Central Limit Theorem. The normal curves show the probability density of the sample mean, highlighting spread and separation between algorithm performances.
To assess the robustness of algorithm rankings, we conducted a bootstrap analysis using 2,000 resampled test sets. Specifically, each bootstrap sample was created by resampling the test cases with replacement, keeping the sample size unchanged. Fig. 8(a) visualises the resulting rankings using a blob plot. The size of each bubble is proportional to the relative frequency of the corresponding ranks obtained across bootstrap samples. The median rank for each algorithm is denoted by a black cross. Notably, all algorithms exhibit perfect ranking consistency: each algorithm achieves the same rank in all 2,000 bootstrap samples, resulting in a single bubble per algorithm with 100% frequency. This suggests that performance differences among the algorithms are highly stable under resampling. Fig. 8(b) shows the estimated sampling distributions of the mean final score for each algorithm [55]. Each distribution is modeled as a Gaussian (normal) curve, where the center of the curve corresponds to the empirical mean of the final score for that algorithm, calculated across all test cases. The spread of the curve is determined by the standard error of the mean, computed as the sample standard deviation divided by the square root of the number of test cases. The tightness and separation of these curves reflect the consistency and distinguishability of algorithm performance.
5.3.2 Team-wise Performance Comparison Across Scan Patterns
Figs. 9 and 10 present the normalised scores and raw error metrics, respectively, for each team across various scan patterns. Local metrics (LPE, LLE) are lowest for straight line shape scans across all teams and highest for S shape scans. GPE and GLE (global metrics) are more consistent across shapes but still slightly higher for S shape scans, especially for methods like MoGLo-Net and PLPPI. This indicates that straight line shape scans are more tractable for the evaluated methods, and S shape scans introduce greater accumulated drift and lower global consistency due to their complex trajectories. Overall, linear trajectories such as those in straight line shape scans are handled more robustly by current methods, while more complex paths, especially S shape scans, lead to significant degradation in both global and local performance. These findings highlight the importance of evaluating reconstruction robustness under diverse motion patterns and reinforce the need for algorithms that generalise effectively across varied scanning conditions. Probe orientation also influences performance: all methods show higher LPE and LLE in parallel scans, while GPE and GLE are comparable across orientations over all methods but MoGLo-Net and PLPPI. This indicates that the parallel orientation may induce more accumulated drift and local misalignment, possibly due to less frame-to-frame overlap. These results highlight the need for methods that are not only scale-robust but also less sensitive to orientations, capable of maintaining accuracy under diverse scanning configurations. Across most methods, proximal-to-distal scans result in slightly higher metric errors compared to distal-to-proximal scans. In contrast, the influence of arm side (left vs. right) appears minimal, with comparable performance observed across both groups. Overall, these trends reveal the impact of scan shape, orientation, and direction on method robustness, and highlight the importance of developing reconstruction algorithms that generalise well across diverse scanning conditions. Performance scores (FS, GS, LS, PS, LMS) remain consistent across different scan patterns, as they are normalised metrics tend to be independent of absolute values.
5.3.3 Performance Analysis Across Scan patterns (Pooled Over All Teams)
Given that scan length (SL) is a critical factor affecting reconstruction performance, primarily due to the cumulative nature inherent in freehand ultrasound reconstruction, this section presents a performance analysis across all methods with respect to both scan patterns and scan length. The observed relationship between performance and scan patterns is consistent with the findings reported in Section 5.3.2.
Figs. 11 and 12 show the distribution of performance scores and error metrics across all methods for each scan, with scans ordered by increasing scan length. The trend suggests that, in general, all error metrics tend to increase with scan length, indicating a potential deterioration in both global trajectory accuracy and local alignment. However, this pattern is not strictly consistent across all cases, and some scans show minimal or no significant degradation. This overall tendency indicates that longer sequences may lead to greater accumulated drift, which affects performance at both global and local spatial scales. Performance scores (FS, GS, LS, PS, LMS) remain relatively stable across scan lengths, as they are normalised metrics designed to reduce sensitivity to overall error magnitude.
Fig. 13 presents the distribution of performance scores and error metrics across all methods for each subject, with subjects ordered by increasing scan length. The performance scores remain relatively stable across subjects. The error metrics demonstrate a positive correlation with scan length, increasing in subjects with longer scans.
Figs. 14 and 15 illustrate the distribution of performance scores and error metrics across individual scans, grouped by three scanning protocols: Straight line shape, C shape, and S shape. Within each protocol, scans are ordered by increasing scan length, enabling assessment of both protocol-specific and length-dependent trends. Across all three protocols, the normalised performance scores (FS, GS, LS, PS, LMS) generally remain within a consistent range, as expected. For the error metrics (GPE, GLE, LPE, LLE), global metrics (GPE, GLE) increase progressively with scan length in all scanning protocols. Local errors (LPE and LLE) show stable trend, suggesting that local alignment is comparatively less sensitive than global consistency. Comparing across scanning protocols, the straight line shape scans yield the lowest and most stable error values across most metrics. In contrast, the S shape scans are associated with the largest median scores and highest variability. C shape scans fall between these extremes, showing moderate errors and variability. Notably, for scans of comparable length, S shape sequences still perform worse, suggesting that both scan length and path complexity jointly negatively impact method effectiveness.
Figs. 16 and 17 present the distribution of performance scores and error metrics across all evaluated methods for scans categorised by probe orientation, parallel and perpendicular, and ordered by increasing scan length within each group. This configuration allows analysis of both orientation-dependent effects and scan length sensitivity. Performance scores are stable. Both global errors and local errors increase with scan length across both probe orientations. However, the increase is more noticeable in the parallel group, where both the magnitude and variability of errors are significantly higher. Overall, these findings demonstrate that both scan length and probe orientation significantly affect reconstruction performance. While performance degrades with increasing scan length in both settings, scans acquired in a parallel orientation exhibit higher global and local errors, along with greater performance variability.
Figs. 18 and 19 show the distribution of performance scores and error metrics across all methods for individual scans, categorised by the scanned arm (left or right) and arranged in ascending order of scan length. Across both groups, normalised scores (FS, GS, LS, PS, LMS) remain within a consistent range overall. Error metrics (GPE, GLE, LPE, LLE) consistently increase with scan length across both left and right arm scans. What is more, the error metrics are generally consistent between left and right arm scans, suggesting that the arm being scanned does not systematically bias metric magnitudes across methods compared to scan length.
Figure 9: Normalised scores for each team across various scan patterns: straight-line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal and proximal-to-distal.
Figure 10: Raw error metrics for each team across various scan patterns: straight-line shape, C shape and S shape; parallel and perpendicular; left arm and right arm; distal-to-proximal and proximal-to-distal.
Figure 11: Performance scores (FS, GS, LS, PS, LMS) for individual scans across all methods, arranged in ascending order of scan length (SL). Higher scores indicate better performance. Scan length is computed as the cumulative distance between the four corner points of consecutive frames.
Figure 12: Error metrics (GPE, GLE, LPE, LLE) for individual scans across all methods. All metrics are in millimeters (mm). Scans are sorted in ascending order of scan length (SL).
Figure 13: (a) Performance scores for individual subjects across all methods, arranged in ascending order of scan length. Higher scores indicate better performance. (b) Error metrics for individual subjects across all methods. Lower error indicates better performance. All metrics are in millimeters (mm). Scans are sorted in ascending order of scan length.
Figs. 20 and 21 explore the influence of scanning direction on performance, comparing scans acquired in a distal-to-proximal (DtP) versus proximal-to-distal (PtD) direction. Both global and local errors increase progressively with scan length in both directions, but with higher magnitudes and greater variability observed in the PtD group. In particular, GPE and GLE values are noticeably elevated in longer PtD scans, indicating more noticeable drift and loss of global alignment. Local metrics (LPE and LLE) also follow this trend but remain comparatively bounded, supporting the idea that global performance is more sensitive to scanning strategy and sequence length. These findings suggest that scanning direction influences the spatial continuity and reconstruction performance of sequential frames. The consistent increase in both global and local errors with scan length across both directions underscores the well-established challenge of long-sequence reconstruction. However, the amplified degradation in PtD scans indicates that the choice of scanning direction is a non-negligible factor in performance, especially for longer or more complex trajectories. Overall, these results highlight the importance of considering scanning direction as a variable in both evaluation and design of spatial tracking reconstruction systems. Future research may investigate direction-sensitive strategies for drift correction to improve robustness across scan patterns.
Figure 14: Distribution of normalised performance scores for individual scans grouped by scanning protocol: Straight line shape, C shape, and S shape. Scans are ordered by increasing scan length within each group.
The trends observed across figures above consistently demonstrate that both global and local error metrics increase with scan length, regardless of scan patterns (e.g., protocol, arm, or orientation). This suggests a strong dependence of metric magnitude on scan length. To validate this observation quantitatively, the correlation between scan length and each metric is assessed using the Pearson correlation coefficient ($r$). For most metrics, $r$ values range from 0.3 to 0.5, indicating moderate positive correlations between scan length and error magnitude. Notably, the four subject-level metrics exhibit stronger correlations, with $r$ values between 0.58 and 0.78, suggesting a substantial association between scan length and performance degradation at the subject level. In all cases, the correlations are statistically significant ($p<0.05$).
5.3.4 Qualitative Results
Fig. 22 presents the qualitative results of the participating teams, illustrating scans corresponding to the best, worst, and median performance based on GPE (subfigures a-c) and LPE (subfigures d-f) metric errors, while accounting for variability across scan patterns. To ensure representative coverage and avoid redundant selection of the same scan pattern, the scan with the next-lowest (or next-highest / next-median) error was selected when appropriate. To show the quantitative difference, the corresponding numerical metric values are provided. Fig. 22 illustrates that accurate local predictions do not necessarily ensure accurate global reconstructions. Additionally, the error magnitude varies across different scans.
Figure 15: Distribution of error metrics for individual scans grouped by scanning protocol: Straight line shape, C shape, and S shape. Scans are ordered by increasing scan length within each group.
Figure 16: Distribution of performance scores (FS, GS, LS, PS, LMS) for individual scans across all methods, categorised by probe orientation (parallel and perpendicular) and ordered by increasing scan length within each group.
Figure 17: Distribution of error metrics (GPE, GLE, LEP, LLE) for individual scans across all methods, categorised by probe orientation (parallel and perpendicular) and ordered by increasing scan length within each group.
Figure 18: Distribution of individual scores across all methods for each scan, categorised by scanned arm (left arm and right arm), and presented in ascending order of scan length.
Figure 19: Distribution of individual raw metric values across all methods for each scan, categorised by scanned arm (left arm and right arm), and presented in ascending order of scan length.
Figure 20: Distribution of individual scores across all methods for each scan, categorised by scanning direction (distal-to-proximal and proximal-to-distal), and presented in ascending order of scan length.
Figure 21: Distribution of individual raw metric values across all methods for each scan, categorised by scanning direction (distal-to-proximal and proximal-to-distal), and presented in ascending order of scan length.
5.3.5 Further Analysis
Across all evaluated methods, a primary limitation is the sensitivity to scan length, as reflected by the correlation between scan length and error magnitude. This suggests that most approaches struggle to maintain lower global error over long sequences, likely due to cumulative drift or limited global context modeling. Furthermore, performance degradation is particularly prominent in geometrically complex trajectories (e.g., S shape scans), indicating an overall lack of robustness to scanning path variability.
While all teams share these global limitations, individual methods demonstrate distinct advantages. Some teams excel in local consistency, achieving low LPE and LLE across a wide range of scans, indicating effective frame-to-frame alignment strategies. Others achieve stronger performance on long scans, suggesting better global modeling. However, no team consistently outperform others across all settings. This lack of generalisation highlights the current trade-offs between local precision and global robustness.
Notably, the positive impact of ensemble learning and pretraining highlights the value of integrating prior visual knowledge, which could be considered when clinically deployed. Taken together, these results suggest that future development should prioritise hybrid approaches that combine local precision, temporal consistency, and robust drift correction. Additionally, performance across scan patterns emphasises the need for improving adaptability in clinical scanning environments.
6 Discussions
While the current Challenge setup offers a comprehensive evaluation of freehand 3D ultrasound reconstruction methods, several limitations remain and highlight opportunities for improvement in future iterations.
Reducing barriers to participation. A key priority for future editions of the TUS-REC Challenge is to lower the technical entry barrier, thereby enabling broader participation from research groups across computer vision, robotics, and medical imaging domains, including those without prior experience in freehand ultrasound reconstruction. Currently, the prerequisite knowledge of ultrasound imaging principles, spatial calibration procedures, and coordinate systems may discourage otherwise capable participants. By providing detailed documentations, the Challenge can broaden accessibility and encourage participation from a more diverse range of research communities. Notably, with existing effort provided by this Challenge paper, this barrier has already been significantly lowered, laying a strong foundation for continued growth and engagement.
Wider anatomical areas and clinical applications. Another limitation of the present Challenge lies in the anatomical scope of the dataset, which is restricted to forearm scans. While the forearm offers a tractable and clinically relevant use case, it presents relatively constrained geometry and motion characteristics. Therefore, it may not fully capture the broader spectrum of challenges encountered in other anatomical regions. This may limit the generalisability of method performance observed in the current setting.
Score weighting and ranking methodologies. In TUS-REC2024 Challenge, a min-max normalisation strategy is applied at the scan level, scaling performance scores into a fixed range of $[0,1]$. While this approach helps reduce the influence of extreme values, it introduces several limitations. For example, it can over-amplify marginal differences in performance when overall variability is low, leading to exaggerated score separation between similarly performing methods. Additionally, the presence of a poorly performing team can artificially boost the normalised scores of others with only slightly better performance. These effects may distort the overall fairness and interpretability of rankings. It is important to acknowledge that all normalisation strategies inherently involve trade-offs, and no single method is universally optimal across all evaluation contexts. Min-max normalisation can suppress the impact of outliers but may exaggerate differences when overall performance is similar. In contrast, z-score normalisation mitigates sensitivity to extreme values but may obscure meaningful differences when the distribution of scores is non-Gaussian. Ultimately, each method emphasises different aspects of relative performance, and the choice of normalisation inevitably shapes the interpretation of results. While careful selection and justification of normalisation methods can improve transparency, there is no definitive solution that fully eliminates bias or distortion in score scaling. Therefore, normalisation should be viewed as a practical necessity with acknowledged limitations, rather than a universally fair metric transformation.
Risk of data leakage. As the Challenge allows multiple submission attempts and uses a fixed test set, there is a risk that participating methods may become inadvertently overfitted to the test data over time. This is especially relevant when teams refine their models based on feedback from repeated evaluations, potentially optimising for the specifics of the test distribution rather than generalisable performance. Such overfitting can undermine the fairness and validity of the ranking results. To mitigate this issue, future editions of the TUS-REC Challenge should consider introducing additional unseen test data in later evaluation phases. This could include a hold-out set only revealed after the final submission deadline or a progressive test set release strategy. Incorporating fresh data would better evaluate the generalisation ability of submitted methods and reduce the likelihood of overfitting to a static benchmark. Moreover, it would more closely reflect real-world deployment conditions, where models must perform reliably on previously unseen patients and scanning conditions.
7 Conclusion
Trackerless 3D freehand ultrasound reconstruction represents a critical advancement in enabling cost-effective, portable, and workflow-friendly 3D imaging solutions for diverse clinical settings. By eliminating the need for external tracking hardware, these methods promise improved accessibility in point-of-care diagnostics and interventional guidance. However, this paradigm shift also introduces new algorithmic challenges in accurate motion estimation under unconstrained probe motion.
TUS-REC2024 Challenge represents a major step forward in benchmarking the current state of the art for trackerless 3D freehand ultrasound reconstruction. With the largest publicly available dataset for this task and participation from leading international research teams, it has successfully enabled a comparative evaluation of modern methods under a standardised test framework. The dataset, containing over two thousand of scans across multiple subjects, provides an invaluable resource for the community and will continue to support method development and reproducibility beyond the Challenge.
The submitted methods reflect a rich diversity of algorithmic strategies, including spatial and temporal modeling, data augmentation, and fusion architectures. These contributions offer a strong foundation for future research. Despite notable advancements, the task is not yet solved to a clinically satisfactory degree. Sensitivity to scan length and scanning protocol reveals that generalisation remains a challenge, and further work is needed to bridge the gap between experimental performance and clinical deployment.
The Challenge website and infrastructure will continue to be available beyond the official competition period, welcoming post-deadline submissions from the research community. It is intended to serve as a long-term benchmark for trackerless freehand ultrasound reconstruction, enabling ongoing method development, reproducibility studies, and performance comparisons as the field advances.
In summary, TUS-REC2024 Challenge provides not only a rigorous benchmark for current methods but also a catalyst for methodological advancement and clinical translation. Its biomedical and technical impact lies in establishing a shared framework to accelerate the development of practical, high-performance trackerless freehand ultrasound systems.
CRediT authorship contribution statement
Qi Li: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review & editing, Challenge organisation. Shaheer U. Saeed: Conceptualization, Resources, Validation, Writing – review & editing, Challenge organisation. Yuliang Huang: Conceptualization, Investigation, Methodology, Resources, Software, Writing – review & editing, Challenge organisation. Mingyuan Luo: Investigation, Methodology, Software, Writing – review & editing. Zhongnuo Yan: Investigation, Methodology, Software, Writing – review & editing. Jiongquan Chen: Investigation, Methodology, Software, Writing – review & editing. Xin Yang: Investigation, Methodology, Software, Writing – review & editing. Dong Ni: Investigation, Methodology, Software, Writing – review & editing. Nektarios Winter: Investigation, Methodology, Software, Writing – review & editing. Phuc Nguyen: Investigation, Methodology, Software, Writing – review & editing. Lucas Steinberger: Investigation, Methodology, Software, Writing – review & editing. Caelan Haney: Investigation, Methodology, Software, Writing – review & editing. Yuan Zhao: Investigation, Methodology, Software, Writing – review & editing. Mingjie Jiang: Investigation, Methodology, Software, Writing – review & editing. Bowen Ren: Investigation, Methodology, Software, Writing – review & editing. SiYeoul Lee: Investigation, Methodology, Software, Visualization, Writing – review & editing. SeonHo Kim: Investigation, Methodology, Software, Visualization, Writing – review & editing. Minkyung Seo: Investigation, Methodology, Software, Visualization, Writing – review & editing. MinWoo Kim: Investigation, Methodology, Software, Visualization, Writing – review & editing. Yimeng Dou: Investigation, Methodology, Software, Writing – review & editing. Zhiwei Zhang: Investigation, Methodology, Software, Writing – review & editing. Yin Li: Investigation, Methodology, Software, Writing – review & editing. Tomy Varghese: Investigation, Methodology, Software, Writing – review & editing. Dean C. Barratt: Funding acquisition, Project administration, Writing – review & editing, Challenge organisation. Matthew J. Clarkson: Funding acquisition, Project administration, Writing – review & editing, Challenge organisation. Tom Vercauteren: Project administration, Supervision, Writing – review & editing, Challenge organisation. Yipeng Hu: Conceptualization, Funding acquisition, Project administration, Resources, Supervision, Writing – review & editing, Challenge organisation.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The training and validation datasets used in TUS-REC2024 Challenge are publicly available at the following repositories: https://doi.org/10.5281/zenodo.11178508, https://doi.org/10.5281/zenodo.11180794, https://doi.org/10.5281/zenodo.11355499, and https://doi.org/10.5281/zenodo.12752245. The baseline model code has been released and can be accessed via GitHub at https://github.com/QiLi111/tus-rec-challenge_baseline. Participants are encouraged to release their code voluntarily. All publicly released code related to the Challenge will be listed on the official Challenge website: https://github-pages.ucl.ac.uk/tus-rec-challenge/TUS-REC2024/.
Acknowledgement
This work was supported by the EPSRC [EP/T029404/1], a Royal Academy of Engineering/Medtronic Research Chair [RCSRF1819\7\734] (TV), Wellcome/EPSRC Centre for Interventional and Surgical Sciences [203145Z/16/Z], and the International Alliance for Cancer Early Detection, an alliance between Cancer Research UK [C28070/A30912; C73666/A31378], Canary Center at Stanford University, the University of Cambridge, OHSU Knight Cancer Institute, University College London and the University of Manchester. TV is co-founder and shareholder of Hypervision Surgical. Qi Li was supported by the University College London Overseas and Graduate Research Scholarships. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Participating teams acknowledge support from the following funding agencies: MUSIC Lab: National Natural Science Foundation of China (Nos. 12326619, 62171290, 62101343), Science and Technology Planning Project of Guangdong Province (2023A0505020002), and Shenzhen-Hong Kong Joint Research Program (SGDX20201103095613036). ISRU@DKFZ: DKFZ (German Cancer Research Center) Heidelberg; DAAD. AMI-Lab: National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS2021-NR059679); Institute of Information & communications Technology Planning & Evaluation (IITP) under the Artificial Intelligence Convergence Innovation Human Resources Development (IITP-2025-RS-2023-00254177) grant funded by the Korea government (MSIT). UW-Madison Elastography Lab: National Heart, Lung, and Blood Institute, grant number 1R01HL147866; National Science Foundation, grant number CNS 2333491.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to improve language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Figure 22: Trajectories of the four corner points from the ground truth (green lines) and model predictions (red lines) on selected scans. (a): straight line shape scan on the left arm along a perpendicular scanning path, from the distal to the proximal direction; (b) S shape scan on the right arm along a perpendicular scanning path, from the proximal to the distal direction; (c): S shape scan on the right arm along a parallel scanning path, from the proximal to the distal direction; (d): C shape scan on the left arm along a perpendicular scanning path, from the distal to the proximal direction; (e): S shape scan on the left arm along a perpendicular scanning path, from the distal to the proximal direction; (f): S shape scan on the right arm along a parallel scanning path, from the distal to the proximal direction.